# The HITRAN2024 methane update

### v1

@@ 2.1.8 NIST, DLR & USTC measurements 5970–6250 cm-1
Figure 4: Comparison of the intensities for the HITRAN2020 (left pane) and HITRAN2024 line lists (right pane) in the 12CH4 tetradecad, including the $2\nu_{3}$ band. The overlap set includes transition intensities determined at NIST (CRDS) and DLR (FTS) in the Doppler-broadened domain.
@@
Agreement with TheoReTS [rey_theorets_2016] is excellent (at the several permille level) for the assigned $2\nu_{3}$ transition intensities, however, the lack of assignments in the experimental line list for many weaker (non-$2\nu_{3}$) transitions makes direct comparison difficult. Additionally, many weaker lines present in TheoReTS have been merged in the experimental line list as they cannot be separated due to finite signal-to-noise or fitting correlation, further complicating direct comparison.
2.1.9 LiPhy DAS measurements of 12CH3D 6204–6394 cm-1
In spite of the small magnitude of the CH3D/CH4 amount ratio, (nominally $5\times 10^{-4}$ at natural abundance), CH3D has a large relative contribution in the region (up to 70 % near 6300 cm-1 at 80 K) [wang_high_2010]) and contributes greatly to the absorption in the 1.58 µm methane transparency window [lu_ch3d_2011]. Nevertheless, the $3\nu_{2}$ band is not isolated but superimposed on a congested spectrum of lines involving other weaker bands which have to be characterized for a proper description of the absorption in the region. As described in [brown_methane_2013], since its 2012 edition, the HITRAN list was extended to the 6204-6510 cm-1 region by including for the 6204–6394 cm-1 interval, the empirical line list constructed in Lu et al. [lu_ch3d_2011] using diode laser absorption spectroscopy (DAS) of highly enriched CH3D at 296 K. This list of about 5500 lines included about 175 rovibrational assignments and more than 2700 empirical values of the lower state energy, $E^{\prime\prime}_{\mathrm{emp}}$, obtained from the ratios of intensities measured at 296 K and 80 K (2T-method).
@@ 2.1.9 LiPhy DAS measurements of 12CH3D 6204–6394 cm-1
2.1.10 PLDL Dual frequency comb 6770–7570 cm-1
@@
Thirteen total spectra were analyzed at five temperatures from 296-1000 K and 18-300 Torr. Simulations in the icosad region with HITRAN2020 show good agreement with the measured spectra at room temperature, but exhibit significant mismatch at higher temperature. This is not surprising given that the empirically based HITRAN2020 icosad line list originated from spectra measured below 280 K (WKLMC [campargue_2013]). The new high-temperature spectra were used to update many HITRAN2020 lower-state energies, line positions and self shifts, self widths, and power-law temperature dependence exponents for self width and shift.
Much of the high-temperature discrepancy between HITRAN2020 and the DCS data was due to methane absorption features that had been given a default value of lower-state energy (such as $E"\textsubscript{HITRAN}=999$ cm-1). Using multispectrum fitting to the high-temperature spectra, 4283 lower-state energies are assigned which had previously been given this default value. Existing lower-state energies were also updated for 92 lines, and the database adds 293 high-temperature lines that improve the database performance above 300 K. The method for these updates is described in detail in [malarich_dual_2021]. Fig.8 in Ref. [malarich_dual_2021] shows the distribution of new lower-state energy values.
@@
2.1.11 IAO & LiPhy CRDS measurements 7606–7919 cm-1
@@ 2.1.12 LiPhy Kitt peak FTS and CRDS measurements 10802–13922 cm-1
At these wavenumbers, the methane spectrum is especially highly congested even at room temperature, and, according to Ref. [campargue_high_2023], part of the absorption shows up as a “quasi-continuum”, which should be added to the absorption lines to fully account for the methane absorption in the region. The “quasi-continuum” represents about one third of the total absorption in the 11000–11400 cm-1 region. Note that the intensity sum of the lines and of the continuum was found to be in good agreement with the intensity predictions of the TheoReTS list [rey_new_2018].
2.2 Updates in HITRAN
The additions and updates from the data described in Section 2.1 is discussed here. As mentioned above, HITRAN2020 was not simply replaced by these lists provided to the HITRAN team. Instead, a series of validation processes involving comparisons with experimental spectra and previous parameters were used.
2.2.1 900–1050 cm-1
The new line list from Nikitin et al. [nikitin_improved_2024] (12CH4, 13CH4, 12CH3D & 13CH3D) completes the lower energy part of the dyad region, adding many lines including those for the lower-abundance 13CH3D isotopologue. Their data were merged directly to the region, replacing the transition wavenumbers for any difference ($\Delta\tilde{\nu}=0$ cm-1) between HITRAN2020 and the new parameters, the lower energy levels for $\Delta E^{\prime\prime}=0.1$ cm-1, and the intensities for $\Delta S_{ij}/S_{ij}=0.1$. About half of the lines in the list served mainly to update the positions and intensities with a few energy levels and assignments, while the other half was entirely new.
2.2.2 2500–3900 cm-1
Figure 7: Residuals obtained at 296 K between experiment from Bertin & Vander Auwersa [bertin_co2_2024] and Voigt calculations from the line lists used for the HITRAN2024 update
The residuals between $S1$, $S2$ and the models calculated using the HITRAN Application Programming Interface (HAPI) with lines from ExoMol (12CH4), MeCaSDa (12CH4 & 13CH4), and HITEMP (12CH4 & 13CH4) were used to compare the quality of each parameter on a line-by-line basis (these residuals are presended for $S2$ in Fig. 7). Despite the good quality of fit, HITEMP was used when all other data were insufficient, since many lines are not fully assigned. By artificially increasing the pressure of the models, we could isolate the missing or strongly shifted lines in every line list, resulting in downward-pointing peaks. Using a combination of derivatives and thresholding, we could automatically find which areas were missing in HITRAN and which line lists could cover them. Because the residuals can be caused by multiple missing lines, every line above a certain intensity in a small window around the region to update was taken from the new line list. We then used the same technique, without the pressure increase, to find and update the weakly shifted lines in HITRAN. In addition to this automated approach, manual selections were made between lists for lines that were incorrectly updated or missed because of the various thresholds.
2.2.3 3900–4900 cm-1
Spectra $S1$ and from Birk et al. [birk_esa_2017] ($S3$), including CRDS spectra not shown here, were used for the octad region. For the same reason as above, $S3$ was used for its window of 3900–4500 cm-1 instead of $S1$.
Figure 8: Residuals obtained at 296 K and 0.17143 hPa between the experiment from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 9: Residuals obtained at 198 K and 0.16941 hPa between the experiment from Ref. [birk_esa_2017] and Voigt calculations from the line lists used for the HITRAN2024 update.
Figure 10: Residuals obtained between experiment from Birk and Voigt calculations from the past and new HITRAN versions
Unlike in the pentad region, several new line lists presented a clear improvement over large portions of the spectrum, as illustrated in Figures 8 and 9. This result alleviated the need for line-by-line analysis, with a few exceptions that were then fixed. In those regions, the new lists completely replaced HITRAN2020, adding new lines and updating existing ones unless no significant changes were made by the conditions detailed in Sec. 2.2.1. The residuals were then analyzed, keeping HITRAN2020 parameters if needed, ensuring that only the best parameters remained. This method was used for the data from Rodina et al. [rodina_improved_2021], between 4100 and 4190 cm-1, improving both CH4, 13CH4, and CH3D. Their line list extended to 4300 cm-1, but that region was previously updated in HITRAN2020 for TROPOMI [hu_operational_2016]. A list provided by Sung et al. for 13CH4 has been added as well between 4000–4700 cm-1, including the region between 4190–4300 cm-1 which replaces some lines, or intensities, incorrectly assigned to the main isotopologues. The list from Nikitin et al. [nikitin_preliminary_2013] improving CH3D between 4000–4554 cm-1 has been added, adding the lines missing in HITRAN2016 [gordon_hitran2016_2017] when the preliminary version of their data was first used. Finally, the empirical list from Toon (atm.161) has been added between 4341–4900 cm-1, improving CH4, 13CH4, and CH3D.
We observed an approximately systematic shift in the transition wavenumber of lines from Daumont et al. [daumont_new_2013]. Applying a constant shift of 0.001 cm-1 appeared to improve the residuals, particularly below 4200 cm-1. The effect is limited above 4375 cm-1 as the addition of ATM list already improved that region appreciably. The final residuals for the latest HITRAN versions are shown in Figure 10, confirming the great improvement above 4375 cm-1, after keeping the region unchanged in the last 3 versions.
2.2.4 4970–5300 cm-1
2.2.5 5900–6530 cm-1
The new list from Ben Fathallah et al. [benfathallah_ch3d_2024] uses their experimental 12CH3D data, and variational calculations if not available, for inclusion in HITRAN2024. More than 5800 new transitions supplement the lower wavenumber range of the previous data. These data comprise 4800 lines providing updates on the positions, intensities, energy levels, as well as new assignments.
2.2.6 6770–7570 cm-1
The new PLDL list [malarich_dual_2021] mostly replaces the 12CH4 transition wavenumbers and some lower energy levels of HITRAN2020 while keeping the old intensities. More than 22,000 transitions were updated using their data, and about 300 new lines were added. Similarly to the DLR/NIST/USTC measurements, the new pressure-induced parameters were used as described in Section 3.
2.2.7 7600–7920 cm-1
2.2.8 10800-14000 cm-1
@@ 2.3 Overview
In total, about 100,000 lines have either been modified or added from the lists described above to the database, increasing the total number of lines from 446,000 in HITRAN2020 to 485,000 now.
@@
The extent of these additions can be visualized in Fig. 12, where we see that 12CH4 lines have been expanded up to 14000 cm-1 and, with the other isotopologues, completed some of the gaps. Additionally, 300 HITRAN2020 lines were removed; however, updated unassigned lines, or those with an incorrect assignment format in HITRAN2020, are counted as removed and added. These constitute the majority of removed lines, because they could not be automatically matched in other line lists. Some of them now have complete or at least partial assignments.
@@
ExoMol [tennyson_2024_2024] 12CH4 550 450 2503 – 4678
MeCaSDa [RICHARD2024109127] 12CH4 70 30 2786 – 4532
13CH4 40 30
Sung et al. [sung_new_2024] 13CH4 5680 5650 4000 – 4700
Nikitin et al., 2013 [nikitin_preliminary_2013] 12CH3D 690 670 4001 – 4537
Rodina et al., 2021 [rodina_improved_2021] 12CH4 4930 4220 4100 – 4190
@@
atm.161 12CH4 22410 23350 4340 – 4899
13CH4 770 790
12CH3D 140 140
Starikova et al., 2024 [starikova_assignment_2024] 13CH4 1640 1640 4970 – 5300
Reed et al., 2025 [reed_multilaboratory_2025] 12CH4 3020 2710 6016 – 6115
Ben Fathallah et al., 2024 [benfathallah_ch3d_2024] 12CH3D 10640 5850 6098 – 6530
Malarich et al., 2021 [malarich_dual_2021] 12CH4 22320 280 6770 – 7570
Nikitin et al., 2024 [nikitin_first_2024] 12CH4 480 480 7629 – 7918
@@
In general, more transition wavenumbers were taken from the new data than line intensities (see Table 2), which stems from our criteria for replacing the existing parameters. From HITRAN2020, about 56,000 of the transition wavenumbers were replaced, with 16,000 for the intensities, 7500 for the energy levels, and 3400 for the assignments. Given the general lower accuracy of the intensities provided, a larger difference between the HITRAN2020 values and the new data was required before updating the value. In most cases, uncertainties were not even reported. The SNR is another contributing factor to consider when most residuals at room temperature and natural isotopic abundance do not always exhibit an obvious improvement from the new lists, especially for weaker lines.
@@ 3.1 USTC, NIST & DLR measurements
For transitions in the $2\nu_{3}$ band of methane, HITRAN2020 only covers the parameters of Voigt profile (VP), i.e., pressure broadening ($\gamma_{0}$) and pressure shifting ($\delta_{0}$). Yet, to achieve greater precision in methane atmospheric simulations, more intricate line shapes and precise parameters are needed. In this regard, we employed the Speed-Dependent Voigt profile (SDVP) corrected for first-order Line-Mixing (LM), abbreviated as LMSDVP, as our base model to analyze the CRDS data. SDVP is a degregation of the pCqSDHC(HT) profile discussed in Ngo et al.[NGO2013JQSRT], which is more accurate than the VP owing to the inclusion of the speed-dependent effect. The involvement of numerous fitting parameters in the HT profile, compounded by the intricate nature of methane spectra, can result in closely numerically correlated spectral line parameters. This can impede the convergence of fitting solutions when analyzing the actual spectra or cause the derived physical parameters to lack physicality. Therefore, we did not use the HT profile. Indeed, the SDVP is adequate for most endeavors of spectral analysis and simulations, but additional treatment is needed due to the high degeneracy of methane states. Thus we introduced the first-order Rosenkranz parameter parameter, $Y_{n}$, to account for line mixing effects [Rosenkranz1975]. From a computational perspective, $Y_{n}$ exhibits a strong correlation with line intensity $S$. Consequently, an independent measurement of line intensity is crucial for acquiring dependable line-mixing parameters. Together with the intensity measurements by NIST and DLR, a global-fit procedure was developed based on the LMSDV model. We also converted the measured nitrogen-broadening coefficients to air-broadening coefficients using the formula $\gamma_{0}^{\text{air}}=0.79\gamma_{0}^{\mathrm{N}_{2}}+0.21\gamma_{0}^{\mathrm{O}_{2}}$, and utilizing polynomial parameters for $\gamma_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. The nitrogen pressure shifting coefficients were similarly scaled as $\delta_{0}^{\text{air}}=0.79\delta_{0}^{\mathrm{N}_{2}}+0.21\delta_{0}^{\mathrm{O}_{2}}$, again utilizing parameters for $\delta_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. Values of $Y_{n}^{\mathrm{N\textsubscript{2}}}$ for the measured CH4–N2 system are converted for CH4–air through multiplication by 0.985, as shown in Tran et al. for the $2\nu_{3}$ band [Tran2010].
Many weaker transitions from 5970 – 6250 cm-1, outside the NIST/DLR/USTC $2\nu_{3}$ overlap set, were updated with Voigt lineshape parameters ($\gamma_{0},\delta_{0}$) recently determined from ambient temperature (nominally 298 K) FTS measurements performed by Birk et al. at DLR. The initial line list was the Nikitin 2017 line list [nikitin_analysis_2017], which was based on the empirical WKLMC line list, updated with the line intensity parameters described in Reed et al. [reed_multilaboratory_2025] and the line positions of Votava et al. [votava_2022], where available. Improved Voigt parameters are reported for approximately 1650 transitions with $S>3\times 10^{-23}$ cm-1/(molec.cm-2) outside the $2\nu_{3}$ band. Lineshape parameters for $2\nu_{3}$ transitions were floated in the fit; however, the values adopted in HITRAN2024 are those determined by USTC above.
@@
The different line profiles used in these studies impact the value of the retrieved coefficients by a small but non-negligible amount. For consistency with the parameters already present in HITRAN, we chose to focus on the Voigt profile. Separating each profile would greatly reduce the amount of data per fit; therefore, we scaled the respective non-Voigt parameters by the average Voigt to non-Voigt ratio for each transition. This approach could only be performed in cases where data were fit with both non-Voigt and Voigt profiles —thus not all studies could be included. To further increase the amount of available data, the CH4–N2 pressure broadening parameters were combined with those of –air, and scaled by a factor of 0.988 taken from Pine & Gabard [pine_speeddependent_2000]. This approach was previously used in other studies [mondelain_measurement_2007] and agrees very well with other measurements (see Fig. 15). Finally, thanks to the weak branch dependence, broadening parameters for the $P$, $Q$, and $R$ branches were combined. Ideally each branch should be separate, however, the broadening measurements in methane are too uncertain for the difference to be reliably quantified.
Using a few assumptions, we can constrain the Padé approximant function to ensure convergence of the fit. First, because pressure broadening is positive-valued, we will force every parameter to be positive for simplicity. This constraint also prevents discontinuities in the curve by removing the possibility of division by zero. Second, based on measurements in various species, not only methane, we can assume that in the $\lim_{J\to\infty}$, $\gamma_{0}$ is equal to zero. For this reason, it is preferable to have a polynomial of greater order in the divisor. We found that a [3/4] function of the form
$\gamma_{0}=\dfrac{a_{0}+a_{1}m+a_{2}m^{2}+a_{3}m^{3}}{1+b_{1}m+b_{2}m^{2}+b_{3}m^{3}+b_{4}m^{4}}$ (1)
@@
The results (Fig. 16) yield only a single function per symmetry. However, given the lack of data at higher quanta and ranking indices, these are the only reliable Padé curves that could be calculated. Nevertheless, they should still provide a good consistent set of parameters for spectroscopic observations. The $A_{1}$ symmetry diverges rapidly from the others. This could simply be a problem with the fit, but again, no assumption at higher quanta should be made, therefore, we decided to keep it. The choice of parameters added to HITRAN2024 was based on the data available. Not every transition used the Padé value. Specifically, if high-precision measurements were available, they were taken instead. This rule holds not only for new data, but those present in HITRAN2020 as well, which were not replaced if the uncertainty is lower. With this approach, we are able to conserve the structures observed in the experiments. The air-broadening half-widths of the $\nu_{3}$ band from HITRAN2020 were replaced using this rule, ensuring a traceable set of data for future updates. Some HITRAN values were found to be unrealistically low or high, possibly caused by typographical errors, and were thus replaced as well. Using this approach, the total number of replaced broadening parameters was 30,000.
@@
With its different symmetry, CH3D must be treated differently, and benefits from a more rigorous separation based on quantum numbers. A better semi-empirical approach developed by Dudaryonok et al. [dudaryonok_semiempirical_2018] seems to produce very good results in the $\nu_{6}$ band. Ignoring the wavenumber dependence, we applied this method for the broadening parameter of every new line added in Sec. 2. The wavenumber dependence of the pressure shift was reintroduced, summing the semi-empirical value to the empirical formula in HITRAN2012 [brown_methane_2013] of $2\times 10^{-6}\times\Delta\nu$, limited to negative values. A similar approach was used where the extremely low or high values of HITRAN2020 were replaced along with the new lines, which only amount to about 70.

### v2

@@ 2.6.1 NIST, DLR & USTC measurements
Figure 8: Comparison of the intensities for the HITRAN2020 (left panel) and HITRAN2024 line lists (right panel) in the 12CH4 tetradecad, including the $2\nu_{3}$ band. The overlap set includes transition intensities determined at NIST (CRDS) and DLR (FTS) in the Doppler-broadened domain.
@@
Agreement with TheoReTS [rey_theorets_2016] is excellent (at the several permille level) for the assigned $2\nu_{3}$ transition intensities, however, the lack of assignments in the experimental line list for many weaker (non-$2\nu_{3}$) transitions makes direct comparison difficult. Limited experimental signal-to-noise or fitting correlation requires the inclusion of many unassigned weak transitions in the experimental linelist to fully model the observed spectra. These transitions likely represent the sum of a number of weaker transitions predicted by THeoReTS.
2.6.2 LIPhy DAS measurements of 12CH3D
In spite of the small magnitude of the CH3D/CH4 amount ratio, (nominally $5\times 10^{-4}$ at natural abundance), CH3D has a large relative contribution to the methane absorption in the region (up to 70 % near 6300 cm-1 at 80 K) [wang_high_2010]) and contributes greatly to the absorption in the 1.58 µm methane transparency window [lu_ch3d_2011]. Nevertheless, the $3\nu_{2}$ band is not isolated but superimposed on a congested spectrum of lines involving other weaker bands which have to be characterized for a proper description of the absorption in the region. As described in [brown_methane_2013], since its 2012 edition, the HITRAN list was extended to the 6204-6510 cm-1 region by including for the 6204–6394 cm-1 interval, the empirical line list constructed in Lu et al. [lu_ch3d_2011] using diode laser absorption spectroscopy (DAS) of highly enriched CH3D at 296 K. This list of about 5500 lines included about 175 rovibrational assignments and more than 2700 empirical values of the lower state energy, $E^{\prime\prime}_{\mathrm{emp}}$, obtained from the ratios of intensities measured at 296 K and 80 K (2T-method).
@@ 2.6.2 LIPhy DAS measurements of 12CH3D
2.6.3 HITRAN update in the 5900–6530 cm-1 region
The new list from Ben Fathallah et al. [benfathallah_ch3d_2024] that uses their experimental 12CH3D, and variational calculations if not available, was included in HITRAN2024. More than 5800 new transitions supplement the lower wavenumber range of the previous data. These data comprise 4800 lines providing updates on the positions, intensities, and energy levels, as well as new assignments.
2.7 6770–7570 cm-1
2.7.1 PLDL Dual frequency comb
@@
Thirteen total spectra were analyzed at five temperatures from 296-1000 K and 18-300 Torr. Simulations in the icosad region with HITRAN2020 show good agreement with the measured spectra at room temperature, but exhibit significant mismatches at higher temperatures. This is not surprising given that the empirically based HITRAN2020 icosad line list originated from spectra measured below 280 K (WKLMC [campargue_2013]). The new high-temperature spectra were used to update many HITRAN2020 lower-state energies, line positions and self shifts, self widths, and power-law temperature dependence exponents for self width and shift.
Much of the high-temperature discrepancy between HITRAN2020 and the DCS data was due to methane absorption features that had been given a default value of lower-state energy (such as $E"\textsubscript{HITRAN}=999$ cm-1). Using multispectrum fitting to the high-temperature spectra, 4283 lower-state energies are assigned, which had previously been given this default value. Existing lower-state energies were also updated for 92 lines, and the database adds 293 high-temperature lines that improve the database performance above 300 K. The method for these updates is described in detail in [malarich_dual_2021]. Fig.8 in Ref. [malarich_dual_2021] shows the distribution of new lower-state energy values.
@@
2.7.2 HITRAN update in the 6770–7570 cm-1 region
The new PLDL list [malarich_dual_2021] mostly replaces the 12CH4 transition wavenumbers and some lower energy levels of HITRAN2020 while keeping the old intensities. More than 22,000 transitions were updated using their data, and about 300 new lines were added. The improved residuals between the measurement and updated database are shown in Fig. 10. Similarly to the DLR/NIST/USTC measurements, the new pressure-induced parameters were used as described in Section 3.
2.8 7600–7920 cm-1
2.8.1 IAO & LIPhy CRDS measurements
@@ 2.9.1 LIPhy Kitt peak FTS and CRDS measurements 10802–13922 cm-1
At these wavenumbers, the methane spectrum is especially highly congested even at room temperature, and, according to Ref. [campargue_high_2023], part of the absorption shows up as a “quasi-continuum”, which should be added to the absorption lines to fully account for the methane absorption in the region. The “quasi-continuum” represents about one third of the total absorption in the 11000–11400 cm-1 region. Note that the intensity sum of the lines and of the continuum was found to be in good agreement with the intensity predictions of the TheoReTS list [rey_new_2018]. In this context, one can use the weak lines from the HITEMP database [10.3847/1538-4365/ab7a1a] (which is largely based on TheoReTS, especially in this region).
2.9.2 HITRAN update in the 10800–14000 cm-1 region
@@ 2.10 Overview
In total, about 125,000 lines have either been modified or added from the lists described above to the database, increasing the total number of lines from 446,000 in HITRAN2020 to 471,000 now.
@@
The extent of these additions can be visualized in Fig. 12, where we see that 12CH4 lines have been expanded up to 14000 cm-1 and, with the other isotopologues, completed some of the gaps. Additionally, 16,000 HITRAN2020 lines were removed; however, updated unassigned lines, or those with an incorrect assignment format in HITRAN2020, are counted as removed and added. These constitute the majority of removed lines, because they could not be automatically matched in other line lists. Some of them now have complete or at least partial assignments.
@@
ExoMol [yurchenko_exomol_2024] 12CH4 240 140 2503 – 4177
MeCaSDa [RICHARD2024109127] 12CH4 70 30 2786 – 3910
13CH4 30 30
Sung et al. [sung_new_2024] 13CH4 4710 4710 4000 – 4700
Nikitin et al., 2013 [nikitin_preliminary_2013] 12CH3D 1530 1510 4001 – 4554
Rodina et al., 2021 [rodina_improved_2021] 12CH4 4930 4230 4100 – 4190
@@
atm.161 [toon_atmospheric_2022] 12CH4 22320 23450 4340 – 4899
13CH4 1710 1700
12CH3D 500 500
Reed et al., 2025 [reed_multilaboratory_2025] 12CH4 7150 7270 6016 – 6115
13CH4 650 660
12CH3D 60 60
Ben Fathallah et al., 2024 [benfathallah_ch3d_2024] 12CH3D 10630 5850 6098 – 6530
Malarich et al., 2021 [malarich_dual_2021] 12CH4 22320 270 6770 – 7570
Nikitin et al., 2024 [nikitin_first_2024] 12CH4 480 480 7629 – 7919
@@
In general, more transition wavenumbers were taken from the new data than line intensities (see Table 2), which stems from our criteria for replacing the existing parameters. From HITRAN2020, about 51,000 of the transition wavenumbers were replaced, with 18,000 for the intensities, 11,000 for the energy levels, and 3300 for the assignments. Given the general lower accuracy of the intensities provided, a larger difference between the HITRAN2020 values and the new data was required before updating the value. In most cases, uncertainties were not even reported. The SNR is another contributing factor to consider when most residuals at room temperature and natural isotopic abundance do not always exhibit an obvious improvement from the new lists, especially for weaker lines.
@@ 3.1 USTC, NIST & DLR measurements
For transitions in the $2\nu_{3}$ band of methane, HITRAN2020 only covers the parameters of Voigt profile (VP), i.e., pressure broadening ($\gamma_{0}$) and pressure shifting ($\delta_{0}$). Yet, to achieve greater precision in atmospheric methane simulations, more intricate line shapes and precise parameters are needed. In this regard, we employed the Speed-Dependent Voigt profile (SDVP) corrected for first-order Line-Mixing (LM), abbreviated as LMSDVP, as our base model to analyze the CRDS data. SDVP is a simplification of the pCqSDHC(HT) profile discussed in Ngo et al.[NGO2013JQSRT], which is more accurate than the VP owing to the inclusion of the speed-dependent effect. The involvement of numerous fitting parameters in the HT profile, compounded by the intricate nature of methane spectra, can result in closely numerically correlated spectral line parameters. This can impede the convergence of fitting solutions when analyzing the actual spectra or cause the derived physical parameters to be less physical. Therefore, we did not use the HT profile. Indeed, the SDVP is adequate for most endeavors of spectral analysis and simulations, but additional treatment is needed due to the high degeneracy of methane states. Thus, we introduced the first-order Rosenkranz parameter, $Y_{n}$, to account for line mixing effects [Rosenkranz1975]. From a computational perspective, $Y_{n}$ exhibits a strong correlation with line intensity $S$. Consequently, an independent measurement of line intensity is crucial for acquiring dependable line-mixing parameters. Together with the intensity measurements by NIST and DLR, a global-fit procedure was developed based on the LMSDV model. We also converted the measured nitrogen-broadening coefficients to air-broadening coefficients using the formula $\gamma_{0}^{\text{air}}=0.79\gamma_{0}^{\mathrm{N}_{2}}+0.21\gamma_{0}^{\mathrm{O}_{2}}$, and utilizing polynomial parameters for $\gamma_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. The nitrogen pressure shifting coefficients were similarly scaled as $\delta_{0}^{\text{air}}=0.79\delta_{0}^{\mathrm{N}_{2}}+0.21\delta_{0}^{\mathrm{O}_{2}}$, again utilizing parameters for $\delta_{0}^{\mathrm{O}_{2}}$ as provided by Lyulin et al. [Lyulin2009JQSRT]. Values of $Y_{n}^{\mathrm{N\textsubscript{2}}}$ for the measured CH4–N2 system are converted for CH4–air through multiplication by 0.985, as shown in Tran et al. for the $2\nu_{3}$ band [Tran2010]. The speed-dependent Voigt parameters were incorporated in HITRAN2024 wherever available.
[…]
@@
The different line profiles used in these studies impact the value of the retrieved coefficients by a small but non-negligible amount. For consistency with the parameters already present in HITRAN, we chose to focus on the Voigt profile. Separating each profile would greatly reduce the amount of data per fit; therefore, we scaled the respective non-Voigt parameters by a constant value to match the Voigt broadening. The constant scaling was calculated by taking the ratio of Voigt to non-Voigt broadening for each transition and averaging the result. This approach could only be performed in cases where transitions were fit with both non-Voigt and Voigt profiles —thus not all studies could be included. To further increase the amount of available data, the CH4–N2 pressure broadening parameters were combined with those of –air, and scaled by a factor of 0.988 taken from Pine & Gabard [pine_speeddependent_2000]. This approach was previously used in other studies [mondelain_measurement_2007] and agrees very well with other measurements (see Fig. 15). Finally, thanks to the weak branch dependence, broadening parameters for the $P$, $Q$, and $R$ branches were combined. Ideally, each branch should be separate; however, the broadening measurements in methane are too uncertain for the difference to be reliably quantified.
Using a few assumptions, we can constrain the Padé approximant function to ensure convergence of the fit. First, because pressure broadening is positive-valued, we will force every parameter to be positive for simplicity. This constraint also prevents discontinuities in the curve by removing the possibility of division by zero. Second, based on measurements in various species, not only methane, we can assume that in the $\lim_{J\to\infty}$, $\gamma_{0}$ is equal to zero. For this reason, it is preferable to have a polynomial of greater order in the divisor. We found that a [2/3] function of the form
$\gamma_{0}=\dfrac{a_{0}+a_{1}m+a_{2}m^{2}}{1+b_{1}m+b_{2}m^{2}+b_{3}m^{3}}$ (1)
@@
Table 3: Padé coefficients of each symmetry fit to the parameters found in literature
$A1$
$A2$
$E$
$F1$
$F2$
$a_{0}$ $4.39417\times 10^{-2}$ $5.69782\times 10^{-2}$ $6.18686\times 10^{-2}$ $5.86367\times 10^{-2}$ $6.12594\times 10^{-2}$
$a_{1}$ $1.60451\times 10^{-2}$ $0.00000$ $0.00000$ $2.02531\times 10^{-4}$ $0.00000$
$a_{2}$ $2.06606\times 10^{-3}$ $1.38168\times 10^{-3}$ $3.11628\times 10^{-4}$ $2.46340\times 10^{-3}$ $1.07334\times 10^{-3}$
$b_{1}$ $0.00000$ $0.00000$ $0.00000$ $0.00000$ $0.00000$
$b_{2}$ $7.54018\times 10^{-2}$ $1.20675\times 10^{-2}$ $6.08344\times 10^{-3}$ $1.27335\times 10^{-2}$ $4.01121\times 10^{-3}$
$b_{3}$ $7.49201\times 10^{-16}$ $1.69122\times 10^{-3}$ $3.30447\times 10^{-4}$ $4.36874\times 10^{-3}$ $1.97354\times 10^{-3}$
The results (Table 3 and Fig. 15 and 16) yield only a single function per symmetry. However, given the lack of data at higher quanta and ranking indices, these are the only reliable Padé curves that could be calculated. Nevertheless, they should still provide a good consistent set of parameters for spectroscopic observations. The $A_{1}$ symmetry diverges rapidly from the others. This could simply be a problem with the fit, but again, no assumption at higher quanta should be made, therefore, we decided to keep it. The choice of parameters added to HITRAN2024 was based on the data available. Not every transition used the Padé value. Specifically, if high-precision measurements were available, they were taken instead. This rule holds not only for new data, but those present in HITRAN2020 as well, which were not replaced if the uncertainty is lower. With this approach, we are able to conserve the structures observed in the experiments. The air-broadening half-widths of the $\nu_{3}$ band from HITRAN2020 were replaced using this rule, ensuring a traceable set of data for future updates. Some previous values were found to be unrealistically low or high, possibly caused by typographical errors, and were thus replaced as well. Using this approach, the total number of replaced broadening parameters was 30,000.
@@
With its different symmetry, CH3D must be treated differently, and benefits from a more rigorous separation based on quantum numbers. The semi-empirical approach developed by Dudaryonok et al. [dudaryonok_semiempirical_2018] seems to produce very good results in the $\nu_{6}$ band. Ignoring the wavenumber dependence, we applied their method for the broadening and temperature-dependence parameters of every new line added in Sec. 2. The wavenumber dependence of the pressure shift was reintroduced, summing the semi-empirical value to the empirical formula in HITRAN2012 [brown_methane_2013] of $2\times 10^{-6}\times\Delta\nu$, limited to negative values. A similar approach was used where the extremely low or high values of HITRAN2020 were replaced along with the new lines, which only amount to about 70.